# Broad Instantaneous Bandwidth Microwave Spectrum Analyzer with a Microfabricated Atomic Vapor Cell

### v1

@@ Abstract
We report on broad instantaneous bandwidth microwave spectrum analysis with hot ${}^{87}$b atoms in a microfabricated vapor cell in a large magnetic field gradient. The sensor is a MEMS atomic vapor cell filled with isotopically pure ${}^{87}$b and $\mathrm{N}_{2}$ buffer gas to localize the motion of the atoms. The microwave signals of interest are coupled through a coplanar waveguide to the cell, inducing spin flip transitions between optically pumped ground states of the atoms. A static magnetic field with large gradient maps the frequency spectrum of the input microwave signals to a position-dependent spin-flip pattern on absorption images of the cell recorded with a laser beam onto a camera. In our proof-of-principle experiment, we demonstrate a microwave spectrum analyzer that has $\approx 1$ GHz instantaneous bandwidth centered at 13.165 GHz, 3 MHz frequency resolution, $2$ kHz refresh rate, and a -23 dBm single-tone microwave power detection limit in 1 s measurement time. A theoretical model is constructed to simulate the image signals by considering the processes of optical pumping, microwave interaction, diffusion of ${}^{87}$b atoms, and laser absorption. We expect to reach more than 25 GHz instantaneous bandwidth in an optimized setup, limited by the applied magnetic field gradient. Our demonstration offers a practical alternative to conventional microwave spectrum analyzers based on electronic heterodyne detection.
@@ I Introduction
Thanks to their ease of use and their potential for low-cost, high-volume production [23], atomic vapor cells containing alkali species are one of the main candidates for commercial quantum sensors [24], which have already been employed in atomic clocks [25], magnetometers [26], gyroscopes [27] and gas sensors [28]. In recent years, atomic vapor cells are also being investigated for detection and imaging of high-frequency electromagnetic fields, using either ground-state atoms to measure the magnetic field component [29, 30, 31, 32, 33] or Rydberg atoms to measure the electric field component [34, 35, 36], from MW up to THz frequencies [37, 38]. The quantum nature of atoms enables measurements of the amplitude or frequency of the electromagnetic field based on well-known atomic properties and fundamental constants, providing intrinsic calibration while at the same time suffering less from crosstalk and electromagnetic interference than conventional antenna systems. While most work has focused on detecting electromagnetic field strength, early experiments have observed frequency information of a sub-MHz-band radio-frequency wave with alkali vapor [39, 40, 41, 42, 43]. Measurements of the frequency spectrum of a broadband MW signal have not yet been demonstrated with atomic vapor cells.
In this work we report on the experimental demonstration and theoretical study of a microwave frequency spectrum analyzer based on ground state alkali atoms in a micro-electro-mechanical systems (MEMS) atomic vapor cell fabricated at wafer-level. Our spectrum analyzer features a broad instantaneous bandwidth and is non-cryogenic, easily-aligned, local oscillator-free, and frequency continuous. After a brief introduction of the operating principle and experimental setup, we show measurement results for a single-tone MW signal and compare it to a theoretical simulation of the atomic spin dynamics and interactions. Various characterization measurements are presented to show the performance of our atomic microwave spectrum analyzer. We demonstrate its capabilities by recording the frequency spectra of a set of frequency-modulated (FM) MW signals and a time-dependent spectrogram of a frequency-swept signal. We discuss the dominant factors influencing several key specifications and present parameters for an optimized setup.
@@ II Principle
Figure 1: Setup and operation of the atomic MW spectrum analyzer. (a) Bottom: An illustration of the key components of the experimental setup. The interior of the MEMS cell containing the atomic vapor is depicted in green, with the long edge positioned directly above the signal trace of the CPW. Blue fan-shaped contours illustrate the $\sigma^{+}$ magnetic component of the MW near field in the $xy$-plane above the CPW. Permanent magnets are positioned so as to generate a large magnetic field with gradient along the $z$-axis. The pump and probe laser beams both propagate in the $x$ direction. The camera displays the atomic spin-flip signal recorded by absorption imaging, reflecting the frequency spectrum of the MW signal on the CPW. Top: Hyperfine energy levels of the ${}^{87}$b ground state as a function of the static magnetic field, with the region across the cell highlighted. (b) Photograph of a MEMS atomic vapor cell employed in the experiment, with a square sensing area of $10\,\mathrm{mm}$ side length and internal thickness $t_{1}=200\,\mu$m, filled with ${}^{87}$b vapor. The thin sidewall with $t_{2}=200\,\mu$m allows for a close approach of the CPW. (c) Hyperfine energy levels of the $5^{2}S_{1/2}$ ground and $5^{2}P_{3/2}$ excited states of ${}^{87}$b at a fixed position in the cell, with the relevant optical and microwave transitions indicated. (d) Experimental pulse sequence (see text).
@@ III Setup and operation
The atomic vapor cell is a miniaturized cell fabricated using wafer-scale MEMS microfabrication techniques [44]. It consists of a patterned silicon layer of thickness $t_{1}=200\,\mu$m, defining the internal thickness of the cell, sandwiched between two glass plates and sealed by anodic bonding. Experiments have been carried out using three different MEMS cells with transverse dimensions of $10\,\mathrm{mm}\times 10\,\mathrm{mm}$ (shown in Fig. 1(b)), $6\,\mathrm{mm}\times 6\,\mathrm{mm}$, and $2\,\mathrm{mm}\times 2\,\mathrm{mm}$, respectively. All cells feature a thin sidewall of $t_{2}=200\,\mu$m thickness, allowing the vapor to closely approach the CPW structure and to couple to the near-field of the guided microwave. The cells are filled using $\mathrm{RbN}_{3}$ UV-decomposition [44] with isotopically selected ${}^{87}\mathrm{Rb}$ atomic vapor and $\mathrm{N}_{2}$ buffer gas to slow down the atomic motion, with buffer gas pressures of 100100100\,100mbar, 110110110\,110mbar, and 127127127\,127mbar, respectively, determined from the collisional broadening and shift of the ${}^{87}$b D2 line in absorption spectroscopy at $T_{\mathrm{fill}}=130^{\circ}\mathrm{C}$ and compared with calculations using the Elecsus software [45, 46]. The diffusion coefficient of ${}^{87}$b in $\mathrm{N}_{2}$ buffer gas is estimated with $D=D_{0}\frac{P_{0}}{P_{\mathrm{N_{2}}}}(\frac{T_{\mathrm{set}}}{T_{0}})^{3/2}$, where $D_{0}=0.1819\,\mathrm{cm}^{2}/\mathrm{s}$ [47] is the diffusion coefficient at $P_{0}=1\,\mathrm{atm}$ and $T_{0}=50^{\circ}\mathrm{C}$, and $P_{\mathrm{N_{2}}}$ is the $\mathrm{N}_{2}$ pressure scaled to $T_{0}$. The oven that heats the cell and stabilizes its temperature at $T_{\mathrm{set}}$ is not shown in Fig. 1(a).
@@
Different MW signals to be analyzed are generated by a signal generator and gated with a switch. After a pre-amplifier, a directional coupler, and a circulator, the MW signal is coupled with net gain $\approx 32\,\mathrm{dB}$ into the CPW via a non-magnetic end-launch SMA connector. At the end of the CPW, it is coupled out of the chip in the same way and sent into a $50\,\Omega$ termination.
@@
The laser beam for optical pumping of the vapor is produced by an external cavity diode laser (ECDL) and pulsed by an acousto-optical modulator (AOM). The probe beam is generated by a distributed feedback laser (DFB) and pulsed by a second AOM in double-pass configuration to increase the extinction ratio. The two beams are expanded to cover the entire cell and propagate along $x$ with the pump beam slightly tilted around $z$ to prevent it from being captured by the camera. A 4f lens system images the cell plane onto a fast scientific complementary metal-oxide-semiconductor (sCMOS) camera. Because the collisional broadening of optical transitions ($\sim\!1.5\,\mathrm{GHz}$) exceeds the magnetic field gradient induced frequency span of the optical transitions across the cell ($\sim 0.6\,\mathrm{GHz}$ for $\sigma_{-}$ and $\sim 0.1\,\mathrm{GHz}$ for $\pi$), the pump and probe beams can address all atoms in the cell for our current parameters. For larger $G$ in future experiments, multiple optical pumping and probing beams in different regions of the cell can be applied.
@@
The hyperfine energy levels of the ${}^{87}$b D2 line at $780$ nm and the optical and microwave transitions involved in the atomic spectrum analyzer operation are shown in Fig. 1(c), for atoms at a fixed position $z_{\omega_{\mathrm{MW}}}$ in the cell. The levels are labeled by the magnetic quantum numbers of the electron spin $m_{J}$ and nuclear spin $m_{I}$ along the magnetic field axis, which are approximately good quantum numbers in the strong static field.
@@
The experimental pulse sequence is shown in Fig. 1(d). It starts with a $10\,\mu$s optical pumping pulse with $150\,\textrm{mW}/\textrm{cm}^{2}$ intensity, which drives the $\sigma_{-}$ optical transition from $\mathop{|\Psi_{1}\rangle}\nolimits$ followed by collisional decay due to the buffer gas. This builds up a steady-state population imbalance between $\mathop{|\Psi_{1}\rangle}\nolimits$ and $\mathop{|\Psi_{8}\rangle}\nolimits$, changing the optical depth OD of the vapor for the probe light by $\Delta\textrm{OD}_{\textrm{pump}}$ compared to the initially unpolarized vapor. Subsequently, the MW signal to be analyzed is pulsed on for the duration $\tau_{\mathrm{MW}}$, during which it drives the $\sigma_{+}$ hyperfine transition between $\mathop{|\Psi_{1}\rangle}\nolimits$ and $\mathop{|\Psi_{8}\rangle}\nolimits$, generating a further change $\Delta\text{OD}_{\textrm{MW}}$ on top of $\Delta\text{OD}_{\textrm{pump}}$. Finally, a linearly polarized probe pulse of duration 1 $\mu$s and intensity $I_{\mathrm{in}}\approx 1.5\,\textrm{mW}/\textrm{cm}^{2}$ is sent to the cell, tuned to resonance with the $\pi$ transition from $\mathop{|\Psi_{1}\rangle}\nolimits$. The probe pulse is attenuated by absorption in the atomic vapor as
@@ IV Results and analysis
Figure 2: (a) Atomic spectrum analyzer image of a single-tone MW signal at $\omega_{\mathrm{MW}}/(2\pi)=13.165~{}\mathrm{GHz}$ and $P_{\mathrm{MW}}=0~{}\mathrm{dBm}$. The image shows $\Delta\textrm{OD}_{\textrm{MW}}$ recorded in the 2 mm cell at $T_{\mathrm{set}}=130.9(1)^{\circ}$ with $\tau_{\textrm{MW}}=1~{}\mu s$ and averaged over $k\!=\!200$ repetitions in 2.2 s total integration time. The image ($369\times 369$ pixels) is rotated by $\sim\!1^{\circ}$ counterclockwise to align the signal with the image columns. (b) Theoretical simulation of the same image re-scaled by a global factor $\beta=0.42$ to match the signal amplitude to the data. (c) Experimental $\Delta\textrm{OD}_{\textrm{MW}}$ at $z=0$ as a function of $y$ in comparison with the simulation. (d) Integrated signal $\Sigma_{y}(\Delta\textrm{OD}_{\textrm{MW}})$ with $y\in\left[0,1.2~{}\textrm{mm}\right]$. The Gaussian fit has a FWHM of $\Delta z_{\textrm{FWHM}}=46(1)~{}\mu\textrm{m}$ corresponding to $\gamma/(2\pi)=4.3(1)~{}\textrm{MHz}$. (e) Signal peak position as a function of the applied MW frequency $\omega_{\mathrm{MW}}$ and linear fit. (f) FWHM of Gaussian fits as in (d) to $195~{}\mu$m wide slices of data as a function of $y$ and exponential fit excluding the data point at $y=0$. The width plateaus at $\gamma/(2\pi)=3.2(2)~{}\textrm{MHz}$.
@@ IV.0.1 Lineshape and spectral resolution
Figure 3: (a) Images of $\Delta\textrm{OD}_{\textrm{MW}}$ for different $\tau_{\mathrm{MW}}$ and other conditions as in Fig. 2(a). (b) Signal width $\gamma$ as a function of $\tau_{\mathrm{MW}}$. Red solid line: fit (see text); yellow dash-dotted line: Fourier limit and intrinsic linewidth $\gamma_{0}$; purple dashed line: diffusion limit; green dotted line: simulation result. Inset: zoom into the region $\tau_{\mathrm{MW}}<1~{}\mu$s.
In Fig. 3(a) we study the dependence of the signal shape on MW pulse duration $\tau_{\mathrm{MW}}$. The length of the signal on the image increases with $\tau_{\mathrm{MW}}$ as the weaker MW field further away from the CPW also begins to change the atomic spin state. At the same time, the signal width increases due to atomic diffusion. The signal width $\gamma$ (plateau values analyzed as in Fig. 2(f)) is shown in Fig. 3(b) as a function of $\tau_{\mathrm{MW}}$. The data is fitted with the heuristic model

### v2

@@ Abstract
We report on broad instantaneous bandwidth microwave spectrum analysis with hot 87Rb atoms in a microfabricated vapor cell in a large magnetic field gradient. The sensor is a MEMS atomic vapor cell filled with isotopically pure 87Rb and $\mathrm{N}_{2}$ buffer gas to localize the motion of the atoms. The microwave signals of interest are coupled through a coplanar waveguide to the cell, inducing spin flip transitions between optically pumped ground states of the atoms. A static magnetic field with large gradient maps the frequency spectrum of the input microwave signals to a position-dependent spin-flip pattern on absorption images of the cell recorded with a laser beam onto a camera. In our proof-of-principle experiment, we demonstrate a microwave spectrum analyzer that has $\approx 1$ GHz instantaneous bandwidth centered around 13 GHz, 3 MHz frequency resolution, $2$ kHz refresh rate, and a -23 dBm single-tone microwave power detection limit in 1 s measurement time. A theoretical model is constructed to simulate the image signals by considering the processes of optical pumping, microwave interaction, diffusion of 87Rb atoms, and laser absorption. We expect to reach more than 25 GHz instantaneous bandwidth in an optimized setup, limited by the applied magnetic field gradient. Our demonstration offers a practical alternative to conventional microwave spectrum analyzers based on electronic heterodyne detection.
@@ I Introduction
Thanks to their ease of use and their potential for low-cost, high-volume production [23], atomic vapor cells containing alkali species are one of the main candidates for commercial quantum sensors [24], which have already been employed in atomic clocks [25], magnetometers [26], gyroscopes [27] and gas sensors [28]. In recent years, atomic vapor cells are also being investigated for detection and imaging of high-frequency electromagnetic fields, using either ground-state atoms to measure the magnetic field component [29, 30, 31, 32, 33] or Rydberg atoms to measure the electric field component [34, 35, 36], from MW up to THz frequencies [37, 38]. The quantum nature of atoms enables measurements of the amplitude or frequency of the electromagnetic field based on well-known atomic properties and fundamental constants, providing intrinsic calibration while at the same time suffering less from crosstalk and electromagnetic interference than conventional antenna systems. While most work has focused on detecting electromagnetic field strength, early experiments have observed frequency information of a sub-MHz-band radio-frequency wave with alkali vapor [39, 40, 41, 42, 43]. Recent work on a Rydberg atom MW receiver demonstrated a real-time bandwidth of 16 MHz [44]. Measurements of the frequency spectrum of a broadband MW signal have not yet been demonstrated with atomic vapor cells.
In this work we report on the experimental demonstration and theoretical study of a microwave frequency spectrum analyzer based on ground state alkali atoms in a micro-electro-mechanical systems (MEMS) atomic vapor cell fabricated at wafer-level. Our spectrum analyzer features a broad instantaneous bandwidth and is non-cryogenic, easy to align, local oscillator-free, frequency-continuous, and only minimally perturbs the incident MW signal. After a brief introduction of the operating principle and experimental setup, we show measurement results for a single-tone MW signal and compare it to a theoretical simulation of the atomic spin dynamics and interactions. Various characterization measurements are presented to show the performance of our atomic microwave spectrum analyzer. We demonstrate its capabilities by recording the frequency spectra of a set of frequency-modulated (FM) MW signals and a time-dependent spectrogram of a frequency-swept signal. We discuss the dominant factors influencing several key specifications and present parameters for an optimized setup.
@@ II Principle
Figure 1: Setup and operation of the atomic MW spectrum analyzer. (a) Bottom: An illustration of the key components of the experimental setup. The interior of the MEMS cell containing the atomic vapor is depicted in green, with the long edge positioned directly above the signal trace of the CPW. Blue fan-shaped contours illustrate the $\sigma^{+}$ magnetic component of the MW near field in the $xy$-plane above the CPW. Permanent magnets are positioned so as to generate a large magnetic field with gradient along the $z$-axis. The pump and probe laser beams both propagate in the $x$ direction. The camera displays the atomic spin-flip signal recorded by absorption imaging, reflecting the frequency spectrum of the MW signal on the CPW. Top: Hyperfine energy levels of the 87Rb ground state as a function of the static magnetic field, with the region across the cell highlighted. (b) Photograph of a MEMS atomic vapor cell employed in the experiment, with a square sensing area of $10\,\mathrm{mm}$ side length and internal thickness $t_{1}=200\,\mu$m, filled with 87Rb vapor. The thin sidewall with $t_{2}=200\,\mu$m allows for a close approach of the CPW. (c) Hyperfine energy levels of the $5^{2}S_{1/2}$ ground and $5^{2}P_{3/2}$ excited states of 87Rb at a fixed position in the cell, with the relevant optical and microwave transitions indicated. (d) Experimental pulse sequence (see text).
@@ III Setup and operation
The atomic vapor cell is a miniaturized cell fabricated using wafer-scale MEMS microfabrication techniques [45]. It consists of a patterned silicon layer of thickness $t_{1}=200\,\mu$m, defining the internal thickness of the cell, sandwiched between two glass plates and sealed by anodic bonding. Experiments have been carried out using three different MEMS cells with transverse dimensions of $10\,\mathrm{mm}\times 10\,\mathrm{mm}$ (shown in Fig. 1(b)), $6\,\mathrm{mm}\times 6\,\mathrm{mm}$, and $2\,\mathrm{mm}\times 2\,\mathrm{mm}$, respectively. All cells feature a thin sidewall of $t_{2}=200\,\mu$m thickness, allowing the vapor to closely approach the CPW structure and to couple to the near-field of the guided microwave. The cells are filled using $\mathrm{RbN}_{3}$ UV-decomposition [45] with isotopically selected ${}^{87}\mathrm{Rb}$ atomic vapor and $\mathrm{N}_{2}$ buffer gas to slow down the atomic motion, with buffer gas pressures of 100100100\,100mbar, 110110110\,110mbar, and 127127127\,127mbar, respectively, determined from the collisional broadening and shift of the 87Rb D2 line in absorption spectroscopy at $T_{\mathrm{fill}}=130^{\circ}\mathrm{C}$ and compared with calculations using the Elecsus software [46, 47]. The diffusion coefficient of 87Rb in $\mathrm{N}_{2}$ buffer gas is estimated with $D=D_{0}\frac{P_{0}}{P_{\mathrm{N_{2}}}}(\frac{T_{\mathrm{set}}}{T_{0}})^{3/2}$, where $D_{0}=0.1819\,\mathrm{cm}^{2}/\mathrm{s}$ [48] is the diffusion coefficient at $P_{0}=1\,\mathrm{atm}$ and $T_{0}=50^{\circ}\mathrm{C}$, and $P_{\mathrm{N_{2}}}$ is the $\mathrm{N}_{2}$ pressure scaled to $T_{0}$. The oven that heats the cell and stabilizes its temperature at $T_{\mathrm{set}}$ is not shown in Fig. 1(a).
@@
Different MW signals to be analyzed are generated by a signal generator (with power $P_{\mathrm{MW}}$) and gated with a switch. After a pre-amplifier, a directional coupler, and a circulator, the MW signal is coupled with net gain $\approx 32\,\mathrm{dB}$ into the CPW via a non-magnetic end-launch SMA connector. At the end of the CPW, it is coupled out of the chip in the same way and sent into a $50\,\Omega$ termination. Alternatively, since the atomic vapor cell does not significantly attenuate the MW signal, the termination can be replaced with a coaxial cable to use the signal for further measurements.
@@
The laser beam for optical pumping of the vapor is produced by an external cavity diode laser (ECDL) and pulsed by an acousto-optical modulator (AOM). The probe beam is generated by a distributed feedback laser (DFB) and pulsed by a second AOM in double-pass configuration to increase the extinction ratio. The two beams are expanded to cover the entire cell and propagate along $x$ with the pump beam slightly tilted around $z$ to prevent it from being captured by the camera. A 4f lens system images the cell plane onto a fast scientific complementary metal-oxide-semiconductor (sCMOS) camera. Because the buffer gas-induced collisional broadening of optical transitions ($\sim\!1.5\,\mathrm{GHz}$) exceeds the magnetic field gradient induced frequency span of the optical transitions across the cell ($\sim 0.6\,\mathrm{GHz}$ for $\sigma_{-}$ and $\sim 0.1\,\mathrm{GHz}$ for $\pi$), the pump and probe beams can address all atoms in the cell for our current parameters. For larger $G$ in future experiments, multiple optical pumping and probing beams in different regions of the cell can be applied.
@@
The hyperfine energy levels of the 87Rb D2 line at $780$ nm and the optical and microwave transitions involved in the atomic spectrum analyzer operation are shown in Fig. 1(c), for atoms at a fixed position $z_{\omega_{\mathrm{MW}}}$ in the cell. The levels are labeled by the magnetic quantum numbers of the electron spin $m_{J}$ and nuclear spin $m_{I}$ along the magnetic field axis, which are approximately good quantum numbers in the strong static field. For the field strengths in our experiment, the $\sigma_{+}$ component of the MW field drives the hyperfine transition between $\mathop{|\Psi_{1}\rangle}\nolimits$ and $\mathop{|\Psi_{8}\rangle}\nolimits$ while the other MW polarization components are off-resonant [31].
@@
The experimental pulse sequence is shown in Fig. 1(d). It starts with a $10\,\mu$s optical pumping pulse with $150\,\textrm{mW}/\textrm{cm}^{2}$ intensity, which drives the $\sigma_{-}$ optical transition from $\mathop{|\Psi_{1}\rangle}\nolimits$ followed by collisional decay due to the buffer gas. This builds up a steady-state population imbalance between $\mathop{|\Psi_{1}\rangle}\nolimits$ and $\mathop{|\Psi_{8}\rangle}\nolimits$, changing the optical depth OD of the vapor for the probe light by $\Delta\textrm{OD}_{\textrm{pump}}$ compared to the initially unpolarized vapor. Subsequently, the MW signal to be analyzed is pulsed on for the duration $\tau_{\mathrm{MW}}$, during which it drives the $\sigma_{+}$ hyperfine transition between $\mathop{|\Psi_{1}\rangle}\nolimits$ and $\mathop{|\Psi_{8}\rangle}\nolimits$, generating a further change $\Delta\text{OD}_{\textrm{MW}}$ on top of $\Delta\text{OD}_{\textrm{pump}}$. For our parameters, $\Delta\text{OD}_{\textrm{MW}}<\Delta\text{OD}_{\textrm{pump}}$ across the cell. Finally, a linearly polarized probe pulse of duration 1 $\mu$s and intensity $I_{\mathrm{in}}\approx 1.5\,\textrm{mW}/\textrm{cm}^{2}$ is sent to the cell, tuned to resonance with the $\pi$ transition from $\mathop{|\Psi_{1}\rangle}\nolimits$. The probe pulse is attenuated by absorption in the atomic vapor as
@@ IV Results and analysis
Figure 2: (a) Atomic spectrum analyzer image of a single-tone MW signal at $\omega_{\mathrm{MW}}/(2\pi)=13.165~{}\mathrm{GHz}$ and $P_{\mathrm{MW}}=0~{}\mathrm{dBm}$. The image shows $\Delta\textrm{OD}_{\textrm{MW}}$ recorded in the 2 mm cell at $T_{\mathrm{set}}=130.9(1)^{\circ}$ with $\tau_{\textrm{MW}}=1~{}\mu s$ and averaged over $k\!=\!200$ repetitions in 2.2 s total integration time. The image ($369\times 369$ pixels) is rotated by $\sim\!1^{\circ}$ counterclockwise to align the signal with the image columns. (b) Theoretical simulation of the same image re-scaled by a global factor $\beta=0.42$ to match the signal amplitude to the data. (c) Experimental $\Delta\textrm{OD}_{\textrm{MW}}$ at $z=0$ as a function of $y$ in comparison with the simulation. (d) Integrated signal $\Sigma_{y}(\Delta\textrm{OD}_{\textrm{MW}})$ with $y\in\left[0,1.2~{}\textrm{mm}\right]$. The Gaussian fit has a FWHM of $\Delta z_{\textrm{FWHM}}=46(1)~{}\mu\textrm{m}$ corresponding to $\gamma/(2\pi)=4.3(1)~{}\textrm{MHz}$. (e) Signal peak position as a function of the applied MW frequency $\omega_{\mathrm{MW}}$ and linear fit. (f) FWHM of Gaussian fits as in (d) to $195~{}\mu$m wide slices of data as a function of $y$ and exponential fit excluding the data point at $y=0$. The width plateaus at $\gamma/(2\pi)=3.2(2)~{}\textrm{MHz}$. Error bars are one standard error (SE) confidence intervals of the fit parameters.
@@ IV.0.1 Lineshape and spectral resolution
Figure 3: (a) Images of $\Delta\textrm{OD}_{\textrm{MW}}$ for different $\tau_{\mathrm{MW}}$ and other conditions as in Fig. 2(a). The signal width $\gamma$ was measured at the indicated positions where power broadening is negligible. (b) Signal width $\gamma$ as a function of $\tau_{\mathrm{MW}}$ (error bars: 1 SE confidence interval). Red solid line: fit (see text); yellow dash-dotted line: Fourier limit and intrinsic linewidth $\gamma_{0}$; purple dashed line: diffusion limit; green dotted line: simulation result. Inset: zoom into the region $\tau_{\mathrm{MW}}<1~{}\mu$s.
@@
In Fig. 3(a) we study the dependence of the signal shape on MW pulse duration $\tau_{\mathrm{MW}}$. The length of the signal on the image increases with $\tau_{\mathrm{MW}}$ as the weaker MW field further away from the CPW also begins to change the atomic spin state. At the same time, the signal width increases due to atomic diffusion. The signal width $\gamma$ (plateau values without power broadening, analyzed as in Fig. 2(f)), is shown in Fig. 3(b) as a function of $\tau_{\mathrm{MW}}$. The data is fitted with the heuristic model